OpenStudy (anonymous):

What is the exact value of the expression Show your work. You can show a square root by writing "sqrt()".

OpenStudy (anonymous):

\[5\sqrt{5}+3\sqrt{20}-6\sqrt{3}\]

OpenStudy (anonymous):

20 = 4* 5

OpenStudy (anonymous):

\[5\sqrt{5} + 6\sqrt{5} - 6\sqrt{3}\]

OpenStudy (anonymous):

nw add

OpenStudy (anonymous):

\[11\sqrt{5} - 6\sqrt{3}\]

OpenStudy (jiteshmeghwal9):

\(5\sqrt{5}+3\times4\sqrt{5}-6\sqrt{3}\)

OpenStudy (jiteshmeghwal9):

\(5\sqrt{5}+12\sqrt{5}-6\sqrt{3}\)

OpenStudy (cwrw238):

3sqrt(20) = 3*sqrt(4)*sqrt(5) = 3*2*sqrt5 = 6 sqrt(5)

OpenStudy (anonymous):

lol..... @jiteshmeghwal9 u r wrong....

OpenStudy (cwrw238):

yahoo 's answer is correct

OpenStudy (anonymous):

Just check ur answer

OpenStudy (cwrw238):

jite made a mistake in simplifying 3 aqrt(20)

OpenStudy (anonymous):

Thanks Guys :)